OpenStudy (anonymous):

Solve the following system: x^2+y^2=9 x^2-y^2=1 The solution set is {( sq. rt of____,____),(sq.rt of____,-_____),(-sq.rt of____,_____),(-sq. rt of___,-_____)}

OpenStudy (btaylor):

If you add vertically, you get 2x^2=10. So \[x= \pm \sqrt{5}\] Then you solve for y in the first equation, and you get y^2=-16. So \[x= \pm 4i\] So your answers are:\[(\pm \sqrt{5} , \pm4i)\]

OpenStudy (anonymous):

but why i?

OpenStudy (btaylor):

because you are taking the square root of a negative number. \[i=\sqrt{-1}\]

OpenStudy (anonymous):

hmmmm

Parth (parthkohli):

You may eliminate by adding both equations

OpenStudy (anonymous):

lets try this one again

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2x^2 = 10 }\) \( \color{Black}{\Rightarrow x^2 = 5 }\) \( \color{Black}{\Rightarrow x = \sqrt5 }\) Now substitute \(x\) back to any of the equations and find what \(y\) is.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5 + y^2 = 9 }\) \( \color{Black}{\Rightarrow y^2 = 4 }\) \( \color{Black}{\Rightarrow \Huge y = 2 }\) Where did that \(i\) come from? o.O

OpenStudy (anonymous):

add both equations together, you get \[2x^2=10\] \[x^2=5\] \[x=\sqrt{5}\] or \[x=-\sqrt{5}\]

Parth (parthkohli):

I always forget the \(\pm \) part ugh

OpenStudy (anonymous):

that's better! a nice real answer

OpenStudy (anonymous):

yeah so there are 4 answer right?

Parth (parthkohli):

lol

Parth (parthkohli):

i like to crack jokes up here. hmm.

OpenStudy (btaylor):

oops...my bad.

OpenStudy (anonymous):

\[\sqrt{5},2\] \[\sqrt{5},-2\] \[-\sqrt{5},2\] \[-\sqrt{5},-2\]

Parth (parthkohli):

That's like all the possible permutations of the + and - signs :P

OpenStudy (anonymous):

y=2.03, -2.03

OpenStudy (anonymous):

to be exact

OpenStudy (anonymous):

thanks guys :D